# Quantum gravitational corrections to Reissner-Nordström black hole thermodynamics and their implications for the weak gravity conjecture

### v3

@@ Abstract
The weak gravity conjecture posits that there should exist at least one state with a charge-to-mass ratio $q/M>1$, ensuring that an extremal charged black hole can still decay. In this paper, we investigate the corrections to the thermodynamics of Reissner-Nordström black holes within the framework of effective field theory. The effective action originates from integrating out massless particles, including gravitons, at the one-loop level. We perform a complete thermodynamic analysis for both non-extremal and extremal black holes. For non-extremal black holes, we identify a relationship between the shift in the charge-to-mass ratio and the thermodynamic stability of the black holes. For extremal black holes, we show that quantum gravity effects naturally lead to the super-extremality of charged black holes, which aligns with known results in the literature.
I Introduction
In Einstein’s general relativity, the extremality bound for a Reissner-Nordström (RN) black hole is given by $\frac{q}{M}\leq 1$ in natural units, where $q$ (implicitly $|q|$) and $M$ are the electric charge and mass of the black hole, respectively. For an extremal charged black hole, if it is assumed to decay into two or more lighter components, then, due to the conservation of energy and charge, one of the decay products must have a charge-to-mass ratio $\frac{q}{M}>1$. If such a super-extremal object does not exist in nature, the decay process would be completely forbidden, and the evolution of black holes would cease at a final stage. Eventually, the universe would be populated by a large number of stable remnants with various charges, which is an unfavorable scenario.
This situation has motivated the so-called weak gravity conjecture (WGC). The WGC suggests that for any $U(1)$ gauge field coupled to gravity, there must exist at least one object with charge $q$ and mass $M$ such that its charge-to-mass ratio exceeds the bound of a large extremal black hole Arkani-Hamed:2006emk . As the name implies, consistent theories must include states where electromagnetic repulsion dominates gravitational attraction. Currently, there are various versions of the WGC, both mild and strong, with different motivations, interpretations, and implications. For an in-depth review of this field, see Ref. Harlow:2022ich .
@@ I Introduction
In this regard, higher-derivative terms should be incorporated into the gravitational effective action as manifestations of quantum gravity effects Boulware:1985wk ; Cardoso:2018ptl ; BarrosoVarela:2023ull . The quantum-corrected effective action can be formally written as
@@
Significant achievements have been gained in recent years Kats:2006xp ; Cremonini:2009ih ; Cremonini:2019wdk ; Cheung:2018cwt ; Cheung:2019cwi ; Noumi:2022ybv . Specifically, four-derivative terms such as $R^{2}$, $R_{\mu\nu}^{2}$ and $R_{\mu\nu\alpha\beta}^{2}$ have been extensively studied, and eq.(2) is utilized to place constraints on the combination of the coefficients of these curvature terms. There is a growing body of evidence in support of the WGC, coming from concrete models constructed from string theories or under certain assumptions. In these studies, the local $\mathcal{O}(R^{2})$ terms are often regarded as making the main contributions to the relevant corrections. However, the loop effects of massless particles including gravitons can dominate over them in the IR regime and provide the leading contributions Harlow:2022ich .
Recently, such loop effects have been considered by Charles Charles:2019qqt and Arkani-Hamed et al. Arkani-Hamed:2021ajd . They found that the loop effects of minimally coupled massless particles indeed satisfy the WGC for non-supersymmetric RN black holes. Initially, unaware of these works, we took a different approach. We observe the existence of a non-local effective action that captures the loop effects of massless particles Donoghue:1994dn ; Donoghue:2014yha ; Barvinsky:1985an ; Calmet:2019eof ; Calmet:2024neu ; Teixeira:2020kew ; qfcs , and its implications for cosmology and various types of black holes have been explored, see e.g. Donoghue:2014yha ; Calmet:2019eof ; Xiao:2021zly . Our motivation, therefore, is to investigate whether such effects can reduce the mass of a RN black hole. Not surprisingly, when taking the extremal limit $T\rightarrow 0$, our results reproduce the findings of Refs.Charles:2019qqt ; Arkani-Hamed:2021ajd for extremal black holes. Nevertheless, our analysis offers a fresh and concise perspective on this problem. Furthermore, the advantage of our formalism is that it enables a detailed thermodynamic analysis of both extremal and non-extremal RN black holes, leading to additional interesting implications.
II One-Loop Effective Action of Quantum Gravity
Thanks to the pioneering works Donoghue:1994dn ; Donoghue:2014yha ; Barvinsky:1985an ; Calmet:2019eof ; Calmet:2024neu ; Teixeira:2020kew ; qfcs , after integrating out the quantum fluctuations of gravitons and other massless particles at one-loop level, the higher-derivative term is given by
@@ II One-Loop Effective Action of Quantum Gravity
Note that the loops involving massless particles generate nonlocal effects, which are reflected by the presence of the nonlocal operator $\ln\left(\frac{\Box}{\mu^{2}}\right)$, where $\Box$ is the d’Alembert operator. By construction, the effective action is renormalization group invariant, that is, it is independent of the renormalization parameter $\mu$ El-Menoufi:2015cqw . For the convenience of manipulation, we will absorb $c_{i}(\mu)$ into its partner $\ln\left(\frac{\Box}{\mu^{2}}\right)$, thereby defining the operator $\ln\left(\frac{\Box}{\mu_{*}^{2}}\right)$. Here, $\mu_{*}$ represents the characteristic energy scale beyond which the effective field-theoretical description becomes invalid. The value of $\mu_{*}$ should be determined by the UV theory of quantum gravity Calmet:2024neu . Without loss of generality, it can be taken to be near the Planck energy. It should be noted that we always work in a regime $r>1/\mu_{*}$, which is below such an energy scale.
We are particularly interested in this effective action for two main reasons. Firstly, the Wilson coefficients $\alpha$, $\beta$, and $\gamma$ of the nonlocal terms are calculable and insensitive to concrete UV theories. This insensitivity represents a model-independent prediction of the effective theory. Since the RN black hole is Ricci-flat, the Wilson coefficient $\alpha$ does not contribute. Below we list the values of $\beta$ and $\gamma$ Donoghue:2014yha ; qfcs :
@@
where $n_{s}$, $n_{f}$, $n_{v}$ denote the numbers of minimally coupled scalars, four-component fermions, additional vectors other than the background $U(1)$ gauge field in the low-energy particle spectrum of nature. The contributions from the fluctuations of the gravitational and Maxwell fields mix together and combine to produce the two numbers $52$ and $398$; see Refs.Bhattacharyya:2012wz ; Charles:2015eha for details.
@@
III Thermodynamic Corrections of the RN Black Hole
@@
Conventionally, the expressions of thermodynamic quantities should be derived from the black hole metric using standard techniques. At the first order in $\epsilon$, the perturbed RN black hole metric can be formally written as $g_{\mu\nu}=\bar{g}_{\mu\nu}+\epsilon g_{\mu\nu}^{(1)}$, where $\bar{g}_{\mu\nu}$ represents the original unperturbed RN metric $ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+r^{2}d\Omega^{2}$, and $f(r)=1-\frac{2M}{r}+\frac{q^{2}}{r^{2}}$. However, determining the exact form of the perturbed metric induced by the nonlocal terms in eq.(3) is far from an easy task. Due to the lack of appropriate techniques to handle the variation $\delta\ln\Box$, the gravitational field equations cannot be derived straightforwardly from the variation analysis of the Lagrangian. Thus, certain assumptions have to be made to obtain the perturbed field equations and the corresponding black hole solution. This has been done in our previous work Xiao:2021zly .
@@
Fortunately, there exists a convenient approach indicating that, at the first order in $\epsilon$, the corrections to the black hole’s thermodynamic quantities from any higher-derivative terms $\epsilon L_{\text{hd}}$ can be directly obtained using the unperturbed metric. This approach was discovered occasionally by Guber et al. Gubser:1998nz and rediscovered and well-explained recently by Reall and Santos Reall:2019sah . For more discussions on its validity and applications, refer to Refs.Caldarelli:1999ar ; Xiao:2022auy ; Xiao:2023two ; Ma:2023qqj . By following this approach, we can deduce the expression of $M_{ext}(q,\epsilon)$ without the need of making unnecessary assumptions.
@@
We now start to evaluate the onshell Euclidean action $I_{\text{tot}}(g_{\mu\nu})=I_{\text{RN}}+\epsilon I_{\text{hd}}$ by substituting $g_{\mu\nu}$ into it. The Euclidean action is directly related to the partition function and free energy of the system. According to the aforementioned approach, when evaluating the Euclidean action at the first order in $\epsilon$, using the perturbed metric $g_{\mu\nu}$ or the unperturbed metric $\bar{g}_{\mu\nu}$ can yield the same result. The reasoning is as follows. The Euclidean action consists of two parts, $I_{\text{RN}}$ and $\epsilon I_{\text{hd}}$, which can be evaluated separately. First, consider the two-derivative part $I_{\text{RN}}\equiv\frac{1}{16\pi}\int_{M}(R-F_{\mu\nu}F^{\mu\nu})$ along with its boundary terms. Since the original RN metric $\bar{g}_{\mu\nu}$ is the extremum of this functional integral, any perturbation around $\bar{g}_{\mu\nu}$ will only cause a correction of $\mathcal{O}(\epsilon^{2})$. That is, $I_{\text{RN}}(g_{\mu\nu})=I_{\text{RN}}(\bar{g}_{\mu\nu})+\mathcal{O}(\epsilon^{2})$ under the condition that $g_{\mu\nu}$ and $\bar{g}_{\mu\nu}$ have the same boundary conditions  [^1] It is understandable that, for a reasonable comparison between $\mathcal{F}[f_{1}(x)]$ and $\mathcal{F}[f_{2}(x)]$ in functional analysis, the two functions $f_{1}(x)$ and $f_{2}(x)$ must satisfy the same boundary conditions.. This analysis also holds when changing the thermodynamic ensemble Cremonini:2019wdk . As a result, when evaluating $I_{\text{RN}}(g_{\mu\nu})$ in an ensemble depending on temperature and charge, the corresponding free energy can be written as
$F_{RN}(T,q)=\frac{3q^{2}+\bar{r}_{h}^{2}}{4\bar{r}_{h}}+\mathcal{O}(\epsilon^{2}).$ (6)
@@
The first term is directly taken from the known formula for the unperturbed RN black hole with the same temperature $T$ and charge $q$, as the boundary conditions must be the same. So $\bar{r}_{h}$ represents the outer horizon radius of the unperturbed black hole, and the expression of $\bar{r}(T,q)$ should be solved from the relation
@@
Next, consider the second part $\epsilon I_{\text{hd}}\equiv\epsilon\int_{\mathcal{M}}L_{\text{hd}}$. It is obvious that $\epsilon I_{\text{hd}}(g_{\mu\nu})=\epsilon I_{\text{hd}}(\bar{g}_{\mu\nu})+\mathcal{O}(\epsilon^{2})$ due to the presence of the factor $\epsilon$. In evaluating $\epsilon I_{\text{hd}}(\bar{g}_{\mu\nu})$, it would be tricky to calculate the terms like $\ln(\frac{\Box}{\mu_{*}^{2}})R_{\mu\nu\alpha\beta}$. However, from dimensional analysis, we can anticipate that its main contribution has the form $-\ln(r^{2}\mu_{*}^{2})R_{\mu\nu\alpha\beta}$, and the explicit calculation can be found in, e.g., Calmet:2019eof ; Delgado:2022pcc ; El-Menoufi:2017kew . Substituting it into $\epsilon L_{hd}$ and calculating the Euclidean integral from the horizon to infinity, we can obtain the expression of $\epsilon I_{\text{hd}}(\bar{g}_{\mu\nu})$.
In total, at the first order in $\epsilon$, the free energy of the perturbed RN black hole is given by
@@
III.1 The non-extremal RN black hole and the shift of charge-to-mass ratio
All thermodynamic quantities can be derived from the free energy. The result is consistent with our previous work Xiao:2021zly , where we explicitly solved the perturbed black hole metric under some reasonable assumptions. For interested reader, the detailed computations of this study are available in the auxiliary Notebook of the arXiv version of the paper.
Specifically, the expressions for the mass and entropy are

### v4

@@ Abstract
In this paper, we investigate the quantum gravitational corrections to the thermodynamical quantities of Reissner-Nordström black holes within the framework of effective field theory. The effective action originates from integrating out massless particles, including gravitons, at the one-loop level. We perform a complete thermodynamic analysis for both non-extremal and extremal black holes, and are mainly concerned about the shift in the charge-to-mass ratio $q/M$ that plays an important role in analyzing the weak gravity conjuecture. For non-extremal black holes, we identify a relationship between the shift in the charge-to-mass ratio and the thermodynamic stability of the black holes. For extremal black holes, we show that quantum gravity effects naturally lead to the super-extremality $q/M>1$ of charged black holes.
journal: Chinese Physics C
[…]
1 Introduction
In Einstein’s general relativity, the extremality bound for a Reissner-Nordström (RN) black hole is given by $\frac{q}{M}\leq 1$ in natural units, where $q$ (implicitly $|q|$) and $M$ are the electric charge and mass of the black hole, respectively. On the other hand, for an extremal charged black hole, postulate that it has the ability to decay into two or more lighter components, then, due to the conservation of energy and charge, one of the decay products must have a charge-to-mass ratio $\frac{q}{M}>1$. If such a super-extremal object doesn’t exist in nature, the decay process would be completely forbidden, and the evolution of black holes would cease at a final stage. Eventually, the universe would be populated by a large number of stable remnants with various charges, which is an unfavorable scenario.
This situation has motivated the so-called weak gravity conjecture (WGC). The WGC suggests that for any $U(1)$ gauge field coupled to gravity, there must exist at least one object with charge $q$ and mass $M$ such that its charge-to-mass ratio exceeds the bound of a large extremal black hole [1]. As the name implies, consistent theories must include states where electromagnetic repulsion dominates gravitational attraction. Currently, there are various versions of the WGC, both mild and strong, with different motivations, interpretations, and implications. For an in-depth review of this field, see Ref. [2].
@@ 1 Introduction
In this regard, higher-derivative terms should be incorporated into the gravitational effective action as manifestations of quantum gravity effects [3, 4, 5]. The quantum-corrected effective action can be formally written as
@@
Significant achievements have been gained in recent years [6, 7, 8, 9, 10, 11]. Specifically, four-derivative terms such as $R^{2}$, $R_{\mu\nu}^{2}$ and $R_{\mu\nu\alpha\beta}^{2}$ have been extensively studied, and eq.(2) is utilized to place constraints on the combination of the coefficients of these curvature terms. There is a growing body of evidence in support of the WGC, coming from concrete models constructed from string theories or under certain assumptions. In these studies, local $\mathcal{O}(R^{2})$ terms are often considered the primary contributors to relevant corrections. However, loop effects of massless particles including gravitons can dominate over these $\mathcal{O}(R^{2})$ terms in the IR regime and provide the leading contributions [2]. This underscores the importance and necessity of considering loop effects in the context of the WGC. Indeed, Refs.[12, 13] investigated loop effects of minimally coupled massless particles and found that the WGC is satisfied for extremal RN black holes.
Given the importance of this problem, we revisit the question from an alternative approach. We observe the existence of a non-local effective action that captures the loop effects of massless particles [14, 15, 16, 17, 18, 19, 20, 21, 22]. Naturally, our motivation is to investigate whether such effects can reduce the mass of a RN black hole at fixed charge $q$. The advantage of our formalism is that it enables us to obtain all thermodynamic quantities for a quantum-corrected RN black hole. We find that the loop effects reduce the mass of RN black holes even far from their extremal limit, specifically, in the regime where the classical black hole parameters satisfy $q\leq\bar{r}_{h}<\sqrt{3}q$, where $\bar{r}_{h}$ denotes the corresponding outer horizon radius. This accommodates the results of Refs.[12, 13] for extremal RN black holes as a special case $\bar{r}_{h}\to q$ of our findings. Notably, $q\leq\bar{r}_{h}<\sqrt{3}q$ also corresponds to the parameter range for thermodynamically stable RN black holes. In addition, our complete thermodynamic analysis interestingly shows that the loop effects simultaneously generate both the super-extremality required by the WGC and the logarithmic entropy corrections, which were previously considered as separate research themes.
The structure of the remaining paper is as follows. In Sec. 2, we present the effective action that describes one-loop corrections to the Einstein-Maxwell theory. In Sec. 3, we compute the quantum corrections to the Euclidean action and free energy of the RN black hole induced by one-loop effects. The subsections there are dedicated to analyzing the thermodynamic corrections for non-extremal and extremal RN black holes respectively. Sec. 4 contains the concluding remarks. Additionally, the Appendix includes a comparison between our effective action and those employed in other related works [12, 13]. For the convenience of interested readers seeking to reproduce our results, all the detailed computations are available in the auxiliary notebook of the arXiv version of this paper.
2 One-Loop Effective Action of Quantum Gravity
Thanks to the pioneering works [14, 15, 16, 17, 18, 19, 20, 21, 22], the quantum corrections to Einstein–Maxwell theory can be organized systematically in an effective field theory expansion. In this framework, one treats the classical metric and gauge field as background fields, integrates out the quantum fluctuations (in particular, the massless modes such as gravitons) in the background-field formalism, and then expands the resulting one-loop effective action in powers of curvatures and derivatives. A key output of this construction is that, besides the local counterterms that renormalize the curvature-squared operators, massless loops unavoidably generate genuinely nonlocal form factors. To the leading order in the curvature expansion, the higher-derivative part can be written as
@@ 2 One-Loop Effective Action of Quantum Gravity
The physical foundation of eq. (3) rests on the standard low-energy effective description of quantum gravity: the local coefficients $c_{i}(\mu)$ encode UV-sensitive matching information and depend on the renormalization prescription, while loops of massless fields induce universal, long-range nonlocalities encoded in operators such as $\ln\left(\frac{\Box}{\mu^{2}}\right)$, where $\Box$ denotes the d’Alembert operator. The computation of one-loop divergences (and hence the relevant Wilson coefficients $\alpha$, $\beta$, and $\gamma$) dates back to the 1970s [14, 15]. The modern effective-field-theory perspective, together with its predictive nonlocal sector, was formalized and systematized in Donoghue’s seminal works [17]. Recently, this framework continues to attract interest and has been applied to quantum gravity phenomenology and black hole physics (see, e.g., [19, 24, 25, 26]). In particular, in cosmological scenarios with a time-dependent background, it has been shown how the effective action yields causal behaviors and manifests nonlocality as a memory effect [18].
By construction, the effective action is renormalization group invariant, namely the explicit $\mu$-dependence of $c_{i}(\mu)$ cancels against that of $\ln\left(\frac{\Box}{\mu^{2}}\right)$ [27]. For later convenience, we will absorb $c_{i}(\mu)$ into its partner $\ln\left(\frac{\Box}{\mu^{2}}\right)$, thereby defining the operator $\ln\left(\frac{\Box}{\mu_{*}^{2}}\right)$. Here, $\mu_{*}$ represents the characteristic energy scale beyond which the effective field-theoretical description breaks down; it should be fixed by the UV completion of quantum gravity [20]. Without loss of generality, we take $\mu_{*}$ to be near the Planck scale. Throughout this paper we work in the regime $r\gg 1/\mu_{*}$, where the effective field theory expansion is under control.
We are particularly interested in this effective action for two main reasons. Firstly, the Wilson coefficients $\alpha$, $\beta$, and $\gamma$ of the nonlocal terms are calculable and insensitive to concrete UV theories. This insensitivity represents a model-independent prediction of the effective theory. Since the RN black hole is Ricci-flat, the Wilson coefficient $\alpha$ does not contribute. Below we list the values of $\beta$ and $\gamma$:
@@
where $n_{s}$, $n_{f}$, $n_{v}$ denote the numbers of minimally coupled scalars, four-component fermions, additional vectors other than the background $U(1)$ gauge field in the low-energy particle spectrum of nature. The contributions from the fluctuations of the gravitational and Maxwell fields mix together and combine to produce the two numbers $52$ and $398$. Note that these numerical coefficients represent one-loop corrections to the Einstein-Maxwell theory (see Refs. [28, 43] for details of the derivation). This explains why they differ slightly from those provided in [18, 19, 20], which correspond to one-loop corrections to Einstein theory.
In principle, in Einstein–Maxwell theory, the effective action (3) may also contain operators involving the Maxwell field strength, such as $R_{\mu\nu\rho\sigma}F^{\mu\nu}F^{\rho\sigma}$ and $F^{4}$. However, using the background-field equations of motion and Bianchi identities, these terms can be eliminated, yielding a purely geometric result [15, 43, 29]. Conversely, such terms can be reintroduced by field redefinitions, which also alter the value of $\beta$. Thus, the coefficients in eq.(4) should be understood as those in a gauge where the Maxwell sector has no logarithmic running coefficients at the one-loop level.
@@
3 Thermodynamic Corrections of the RN Black Hole
@@
Conventionally, the expressions of thermodynamic quantities should be derived from the black hole metric using standard techniques. At the first order in $\epsilon$, the perturbed RN black hole metric can be formally written as
$g_{\mu\nu}=\bar{g}_{\mu\nu}+\epsilon g_{\mu\nu}^{(1)},$ (6)
@@
where $\bar{g}_{\mu\nu}$ represents the original unperturbed RN metric $ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+r^{2}d\Omega^{2}$, and $f(r)=1-\frac{2M}{r}+\frac{q^{2}}{r^{2}}$. However, determining the exact form of the perturbed metric induced by the nonlocal terms in eq.(3) is far from an easy task. Due to the lack of appropriate techniques to handle the variation $\delta\ln\Box$, the gravitational field equations cannot be derived straightforwardly from the variation analysis of the Lagrangian. Thus, certain assumptions have to be made to obtain the perturbed field equations and the corresponding black hole solution. This has been done in our previous work [23].
@@
Anyway, next we will adopt another method to dealing with this problem independent of the techniques used in Ref.[23] [^1] We simply use Ref.[23] for the purpose of cross-checking of the results.. We notice that there exists a convenient approach indicating that, at the first order in $\epsilon$, the corrections to the black hole’s thermodynamic quantities from any higher-derivative terms $\epsilon L_{\text{hd}}$ can be directly obtained using the unperturbed metric $\bar{g}_{\mu\nu}$, without knowing the exact form of $g_{\mu\nu}=\bar{g}_{\mu\nu}+\epsilon g_{\mu\nu}^{(1)}$. This approach was discovered occasionally by Guber et al. [30] and rediscovered and well-explained recently by Reall and Santos [52]. For more discussions on its validity and recent applications, refer to Refs.[31, 32, 33, 34, 35, 36, 37, 38]. By following this approach, we can deduce the expression of $M_{ext}(q,\epsilon)$ without the need of making unnecessary assumptions.
We now start to evaluate the onshell Euclidean action
$I_{\text{tot}}(g_{\mu\nu})=I_{\text{RN}}+\epsilon I_{\text{hd}},$ (7)
by substituting $g_{\mu\nu}$ into it. The Euclidean action is directly related to the partition function and free energy of the system. According to the aforementioned approach, when evaluating the Euclidean action, using the perturbed metric $g_{\mu\nu}$ or the unperturbed metric $\bar{g}_{\mu\nu}$ can yield the same result at the first order in $\epsilon$. The reasoning is as follows.
@@
The onshell Euclidean action (7) comprises two components: $I_{\text{RN}}$ and $\epsilon I_{\text{hd}}$, which can be evaluated separately. First, consider the two-derivative term $I_{\text{RN}}\equiv\frac{1}{16\pi}\int_{M}(R-F_{\mu\nu}F^{\mu\nu})$ along with its boundary term. Since the original RN metric $\bar{g}_{\mu\nu}$ corresponds to the extremum of this functional integral, any perturbation around $\bar{g}_{\mu\nu}$ will only induce a correction of $\mathcal{O}(\epsilon^{2})$. This implies that $I_{\text{RN}}(g_{\mu\nu})=I_{\text{RN}}(\bar{g}_{\mu\nu})$ at $\mathcal{O}(\epsilon)$, provided that $g_{\mu\nu}$ and $\bar{g}_{\mu\nu}$ obey the same boundary conditions [^2] In functional analysis, a meaningful comparison between $\mathcal{F}[f_{1}(x)]$ and $\mathcal{F}[f_{2}(x)]$ requires the functions $f_{1}(x)$ and $f_{2}(x)$ to obey the same boundary conditions. For the present application, this means the perturbed metric $g_{\mu\nu}$ and unperturbed metric $\bar{g}_{\mu\nu}$ must share the same temperature and charge.. This analysis also holds when changing the thermodynamic ensemble [8].
As a result, when evaluating the first part $I_{\text{RN}}(g_{\mu\nu})$ of the Euclidean action (7), one can directly adopt the well-known textbook expression for the Euclidean action $I_{\text{RN}}(\bar{g}_{\mu\nu})$ of the original RN black hole. In an ensemble characterized by temperature $T$ and charge $q$, the corresponding free energy is given by
$F_{\text{RN}}(T,q)=\frac{3q^{2}+\bar{r}_{h}^{2}}{4\bar{r}_{h}},$ (8)
which is directly borrowed from the known formula for the unperturbed RN black hole with the same temperature $T$ and charge $q$. Note that passing from the $(T,\Phi)$ ensemble to the $(T,q)$ ensemble can be strictly realized by a Legendre transformation, i.e., adding a total derivative term $\frac{1}{4\pi}\int_{\mathcal{M}}\nabla_{\mu}(F^{\mu\nu}A_{\nu})$ to the action (1).
@@
In eq.(8), $\bar{r}_{h}$ represents the outer horizon radius of the unperturbed black hole, and the expression of $\bar{r}_{h}(T,q)$ should be solved from the relation
@@
Next, consider the second part $\epsilon I_{\text{hd}}\equiv\epsilon\int_{\mathcal{M}}L_{\text{hd}}$. It is obvious that $\epsilon I_{\text{hd}}(g_{\mu\nu})=\epsilon I_{\text{hd}}(\bar{g}_{\mu\nu})+\mathcal{O}(\epsilon^{2})$ due to the presence of the factor $\epsilon$. Consequently, the unperturbed RN metric $\bar{g}_{\mu\nu}$ can be directly employed for the calculation, since we only need the results at $\mathcal{O}(\epsilon)$. The other problem is that the calculation of the terms like $\ln(\frac{\Box}{\mu_{*}^{2}})R_{\mu\nu\alpha\beta}$ would be tricky. However, progresses have been made in Ref.[19, 44, 45], where explicit calculations show that its main contribution has the form $\ln(r^{2}\mu_{*}^{2})R_{\mu\nu\alpha\beta}$, after appropriate regularization procedure. Substituting it into $\epsilon\int_{\mathcal{M}}L_{hd}$ and calculating the integral from the horizon to infinity, we can obtain the expression of $\epsilon I_{\text{hd}}(\bar{g}_{\mu\nu})$. Besides, for the asymptotical-flat cases, the boundary terms associated with the higher derivative part of the action decay rapidly as $r\rightarrow\infty$, therefore can be neglected in the evaluation of $I_{\text{hd}}$ [52].
In total, at the first order in $\epsilon$, the free energy of the perturbed RN black hole is obtained as
$F(T,q)\!=\!\frac{3q^{2}\!+\!\bar{r}_{h}^{2}}{4\bar{r}_{h}}\!-\!\frac{16\pi\epsilon\left((\beta\!+\!4\gamma)q^{4}\!-\!5\gamma q^{2}{\bar{r}_{h}}^{2}\!+\!5\gamma{\bar{r}_{h}}^{4}\right)\ln\left({\bar{r}_{h}}^{2}\mu_{*}^{2}\right)}{5{\bar{r}_{h}}^{5}}.$
@@
3.1 The non-extremal RN black hole and the shift of charge-to-mass ratio
All thermodynamic quantities can be derived from the free energy $F(T,q)$ of the ensemble depending on $T$ and $q$. The entropy and eletrostatic potential of the system can be calculated from $S=-\frac{\partial F}{\partial T}$ and $\Phi=\frac{\partial F}{\partial q}$, and the mass can be calculated from $M=F+TS$. The results are consistent with our previous work [23], where we explicitly solved the perturbed black hole metric under some reasonable assumptions.
Specifically, the mass and entropy are respectively obtained as
$M\!=\!\frac{q^{2}\!+\!\bar{r}_{h}^{2}}{2\bar{r}_{h}}\!+\!\frac{32\pi\epsilon q^{4}(\beta\!+\!4\gamma)\left(2\bar{r}_{h}^{2}\!-\!q^{2}\right)\ln\left(\bar{r}_{h}^{2}\mu_{*}^{2}\right)}{5\bar{r}_{h}^{5}\left(\bar{r}_{h}^{2}-3q^{2}\right)},$